OpenStudy (anonymous):

What is the missing exponent? (2-5) = 2-15

OpenStudy (kailee1423):

What are the options? 2-5=-3 and 2-15=-13

Nnesha (nnesha):

well it's \[\huge\rm (2^{-5})^{?}=2^{-15}\] lets say missing exponent is x \[\huge\rm (2^{-5})^x=2^{-15}\] now apply the exponent rule \[\huge\rm (x^m)^n =x^{m \times n}\]

OpenStudy (anonymous):

ok thanks

Nnesha (nnesha):

got it ?

Nnesha (nnesha):

what did you get ?